# Fiducial Inference for Random-Effects Calibration Models: Advancing Reliable Quantification in Environmental Analytical Chemistry

### v1

@@ 6 Data Analysis
After assessing the model fit, the test samples with unknown concentrations are now analyzed, leading to the construction of confidence intervals and point estimates for the unknown concentrations. Table 5 presents point estimates calculated using both the fiducial method and the MLE for each concentration. Accompanying these point estimates are 95% confidence intervals derived from the proposed fiducial method in Section 4, as well as those obtained using the bootstrap and Wald methods based on MLE.
@@
Cadmium Data Copper Data
True Point Estimate 95% Confidence Interval True Point Estimate 95% Confidence Interval
Conc. Fiducial MLE Fiducial Bootstrap MLE Conc. Fiducial MLE Fiducial Bootstrap MLE
0 1.40 1.51 (0, 5.6) (0, 5.4) (0, 5.0) 0 0.14 0.20 (0, 2.7) (0, 2.1) (0, 2.1)
20 20.5 19.4 (17, 24) (16, 24) (16, 23) 2 2.6 2.6 (0.6, 5.0) (1.1, 4.7) (0.7, 4.5)
100 103.2 102.8 (91, 107) (96, 110) (96, 110) 10 9.1 8.9 (6.5, 11.5) (6.8, 11.0) (7.0, 10.8)
50 47.3 47.1 (42.6, 51.5) (44.7, 49.8) (44.7, 49.5)
200 197.1 195.9 (181, 208) (189, 205) (189, 203)
It is noteworthy that unlike the fiducial confidence intervals whose coverage probabilities are quite accurate, the bootstrap and Wald MLE methods result in coverage probabilities that can be significantly below the assumed confidence level. Clearly, this has implications in the application of these methodologies. For example, for the copper data, when the true concentration is 50, both the bootstrap and Wald MLE methods yield upper limits of the 95% interval that fall below 50. This demonstrates the potential unreliability of these methods.
@@ 7 Discussion
The fiducial approach developed in this article is unique in that it allows estimates for a single laboratory or multiple laboratories analyzing split samples from the same environmental source to borrow strength from other laboratories for which calibration data using the same analytic methodology are available. The distinct advantage of this approach is that both the point estimate and confidence interval are more precise. The approach presented in this paper addresses a crucial challenge in analytical chemistry, namely the uncertainty in chemical measurements. Uncertainty is frequently underestimated in this field, leading to the erroneous belief that measurement estimates reflect true values. The techniques devised in this study facilitate the estimation of uncertainty related to the actual true concentration upon which these measurements are predicated. Given that measurements are typically collected from various instruments, analyses, and often across multiple laboratories, it is essential to consider these sources of variance when evaluating the overall uncertainty in the analytical process. Therefore, these variances should be factored into calculations that define the bounds for the underlying true concentration. In practical applications, when assessing a measurement against a regulatory threshold limit, or compliance standard (referred to as RT), we can compute a confidence interval for the true unknown concentration $X$. An exceedance is declared only if the lower confidence limit (LCL) exceeds the RT, providing 100$(1-\alpha)\%$ confidence that the RT has been surpassed. Thus, determination of the LCL with the prespecified confidence level is extremely important for comparisons to the RT, and our fiducial approach provides an LCL for this purpose.
@@
Supplementary Material outlines the sequential steps required to generate the fiducial pivot quantities for estimating unknown concentrations, including the procedure for handling new test samples. In addition, the supplement discusses specific cases where traditional methods like MLE and MME become computationally infeasible due to limited sample sizes. It also includes comprehensive simulation results with complete tables and graphical analyses that demonstrate the performance of the fiducial approach covering both interval and point estimation, as well as detailed goodness‐of-fit assessments for the Rocke–Lorenzato calibration model on cadmium and copper datasets.

### v2

@@ 6 Data Analysis
After assessing the model fit, the test samples with unknown concentrations are now analyzed, leading to the construction of confidence intervals and point estimates for the unknown concentrations. Table 6 presents point estimates calculated using both the fiducial method and the MLE for each concentration. Accompanying these point estimates are 95% confidence intervals derived from the proposed fiducial method in Section 4, as well as those obtained using the bootstrap, Wald methods based on MLE and LRT by Zhao et al. (2021).
@@
Cadmium Data
True Point Estimate 95% Confidence Interval
Conc. Fiducial MLE Fiducial Bootstrap MLE LRT
0 1.40 1.51 (0, 5.6) (0, 5.4) (0, 5.0) (0, 5.8)
20 20.5 19.4 (17, 24) (16, 24) (16, 23) (16, 24)
100 103.2 102.8 (91, 107) (96, 110) (96, 110) (95, 112)
Copper Data
True Point Estimate 95% Confidence Interval
Conc. Fiducial MLE Fiducial Bootstrap MLE LRT
0 0.14 0.20 (0, 2.7) (0, 2.1) (0, 2.1) (0, 3)
2 2.6 2.6 (0.6, 5.0) (1.1, 4.7) (0.7, 4.5) (0, 6)
10 9.1 8.9 (6.5, 11.5) (6.8, 11.0) (7.0, 10.8) (6.1, 11.8)
50 47.3 47.1 (42.6, 51.5) (44.7, 49.8) (44.7, 49.5) (38.9, 48.5)
200 197.1 195.9 (181, 208) (189, 205) (189, 203) (186, 204)
Notably, we observed in the simulation studies that unlike fiducial confidence intervals, which exhibit highly accurate coverage probabilities, the bootstrap, Wald MLE, and Zhao’s LRT methods often yield coverage probabilities that fall significantly below the intended confidence level. Clearly, this has implications for the application of these methodologies. For example, for the copper data, when the true concentration is 50, both the bootstrap and Wald MLE methods yield upper limits of the 95% interval that fall below 50. This demonstrates the potential unreliability of these methods.
@@ 7 Discussion
The fiducial approach developed in this article is unique in that it allows estimates for a single laboratory or multiple laboratories analyzing split samples from the same environmental source to borrow strength from other laboratories for which calibration data using the same analytic methodology are available. The distinct advantage of this approach is that both the point estimate and confidence interval are more precise, and computationally it is easy to implement. The approach presented in this paper addresses a crucial challenge in analytical chemistry, namely the uncertainty in chemical measurements. Uncertainty is frequently underestimated in this field, leading to the erroneous belief that measurement estimates reflect true values. The techniques devised in this study facilitate the estimation of uncertainty related to the actual true concentration upon which these measurements are predicated. Given that measurements are typically collected from various instruments or analyses, and often across multiple laboratories, it is essential to consider these sources of variability when evaluating the overall uncertainty in the analytical process. Therefore, these variances should be factored into calculations that define the bounds for the underlying true concentration. In practical applications, when assessing a measurement against a regulatory threshold limit, or compliance standard (referred to as RT), we can compute a confidence interval for the true unknown concentration $X$. An exceedance is declared only if the lower confidence limit (LCL) exceeds the RT, providing 100$(1-\alpha)\%$ confidence that the RT has been surpassed. Thus, determination of the LCL with the prespecified confidence level is extremely important for comparisons to the RT, and our fiducial approach provides an LCL for this purpose.
@@
Supplementary Material discusses specific cases where traditional methods like MLE and MME become computationally infeasible due to limited sample sizes. It also includes comprehensive simulation results with complete tables and graphical analyses that demonstrate the performance of the fiducial approach covering both interval and point estimation, as well as detailed goodness‐of-fit assessments for the Rocke–Lorenzato calibration model on cadmium and copper datasets.